# A Simple Derived Categorical Generalization of Ulrich Bundles

### v1

@@ Abstract.
This paper gives a characterization of the Ulrich object similar to the one in [ES03]. As an application, we have provided a new approach to the Eisenbud-Schreyer question by using the notions of the generator of the derived category. We also have given an example of Ulrich objects that are not sheaf by the Yoneda extension.
@@
Derived category, Ulrich bundles, Generators of triangulated categories, Algebraic curves, Bridgeland stability condition
@@ 0. Introduction
Let $X$ be a smooth projective variety over $k=\overline{k}$ and $D^{b}(X)$ be the bounded derived category of coherent sheaves on $X$. A derived category, whose objects are bounded complexes of coherent sheaves, is a triangulated category constructed from an abelian category. The investigation of $D^{b}(X)$ as an invariant of variety began with the work of Bondal and Olrov [BO01], where they showed that if $X$ has (anti-)ample canonical line bundle, then the equivalence $D^{b}(X)\cong D^{b}(Y)$ implies $X\cong Y$.
@@
Meanwhile, recent researches has shown that the derived categories are also valuable to investigate the properties of sheaves. For instance, derived categorical methods are employed in the works of Ohno [Ohn20, Ohn22] and Fukuoka, Hara, and Ishikawa [FHI22, FHI23] to classify certain classes of special bundles.
@@ 0.3. Application to the Eisenbud-Schreyer’s Question
We present a new general proof of the existence of Ulrich bundles on an algebraic curve with an arbitrary embedding, based on the framework of the derived category. Note that the existence of Ulrich line bundle on a curve has already established in [ES09]. To approach the existence problem, our proof is divided into two parts. The first part is the following proposition:
Proposition C (see 4.9).
Let $\mathcal{A}$ be a hereditary abelian category and $\mathcal{D}(\mathcal{A})$ the derived category of $\mathcal{A}$ with arbitrary direct sums. If $G\in\mathcal{D}^{\mathrm{cpt}}(\mathcal{A})$ is a weak generator of $\mathcal{D}^{\mathrm{cpt}}(\mathcal{A})$, then $G$ is also a generator of $\mathcal{D}(\mathcal{A})$.
The second step is showing that $\mathcal{G}^{\prime}$ cannot classically generate $D^{b}(C)$. By combining B and C, we arrive at the following corollary:
Theorem D (see 4.10).
For any smooth projective curve with an embedding $(C,\mathscr{O}_{C}(1))$, there exists an Ulrich object.
@@
The existence problem has been widely studied and there are too many references to list them out here, but we mention some of them:
@@
We believe this method can be extended to more general cases, potentially providing a comprehensive answer to the A.
@@
The following Section 3 discusses the example of Ulrich objects. In particular, we introduce a special type of Ulrich object, called Yoneda type Ulrich object. The application of Ulrich object to A is given in Section 4. Especially, the proof of D is in Section 4.2. Finally, some remarks about the relationship between the Ulrich objects and (divisorial) Bridgeland stability conditions are in Section 5.
@@
The convention and notations used in this paper are listed below:
@@
  • all functors are derived unless otherwise stated. However, $\mathbf{L}$ and $\mathbf{R}$ are occasionally used to emphasize the functor is actually derived.
@@ Acknowledgement.
The aurhor is grateful to his advisor Professor Yasunari Nagai for his continuous encouragement and fruitful discussions. He would also like to thank Professors Wahei Hara and Yuki Mizuno for useful discussions. The author was supported by JST SPRING, Grant Number JPMJSP2128.
@@ 1.1. Preliminaries on Ulrich Bundles
$0\to\mathscr{O}_{\mathbb{P}^{n+1}}(-1)^{\oplus rd}\stackrel{{\scriptstyle M}}{{\longrightarrow}}\mathscr{O}_{\mathbb{P}^{n+1}}^{\oplus rd}\to\mathcal{E}\to 0$
@@ 3.1. Yoneda Type Ulrich Objects
$0\to\mathcal{G}\to E_{k-1}\to\cdots\to E_{0}\to\mathcal{F}\to 0$
@@
[…]
@@ 3.2. Explicit Examples
When the dimension of the variety is 11 11, Ulrich objects are just a direct sum of Ulrich bundles from 2.2. Indeed, in this case, any object $E$ is formal, that is, $E$ can be written as
@@ 3.2.2. Ulrich Objects on Surfaces
Thus, giving an Ulrich object on a surface $S$ is equivalent to giving finitely many Ulrich sheaves $\mathcal{G}^{i}$ and their degree 22 22 extensions.
Moreover, the Ulrichness of the object imposes a strong restriction on its Chern character. To see this, for example, let $X$ be a surface of Picard rank 11 11. Let us write $\operatorname{Pic}(X)=\mathbb{Z}[H]$, $H^{2}=d$, and $K_{X}=i_{X}H$.
@@ 4.1. Generators of the Derived Category
=$\displaystyle=\{E\in\mathcal{D}\mid\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(X,E)=0,{}^{\forall}X\in\Omega\}.$
@@
Proposition 4.5 ([Sta18, Tag 09SR], [Nee92]).
Let $\mathcal{D}$ be a triangulated category with direct sums and $G$ be a compact object of $\mathcal{D}$. Then, the following are equivalent:
  1. (1)
    $G$ is a classical generator for $\mathcal{D}^{\mathrm{cpt}}$ and $\mathcal{D}$ is compactly generated, and
  2. (2)
    $G$ is a generator for $\mathcal{D}$.
4.2. Existence for Curves
In this Section 4.2, we will show D.
Lemma 4.6.
Let $\mathcal{D}$ be a triangulated category with arbitrary direct sums and $G$ a weak generator of $\mathcal{D}^{\mathrm{cpt}}$. Let $f:A\to B$ be a morphism in $D$ such that $C\coloneqq\operatorname{Cone}f\cong\bigoplus_{i}E_{i}$, where $E_{i}\in\mathcal{D}^{\mathrm{cpt}}(X)$. Then,
$f\circ-:\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,A)\stackrel{{\scriptstyle\cong}}{{\to}}\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,B)$
if and only if $f$ is an isomorphism.
@@ 4.2. Existence for Curves
By applying $\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,-)$ to the triangle defined by $f$, the following triangle arises:
$\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,A)\to\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,B)\to\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,C)\cong\bigoplus_{i}\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,E_{i})\stackrel{{\scriptstyle[1]}}{{\to}}.$
Thus, if $f\circ-:\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,A)\stackrel{{\scriptstyle\cong}}{{\to}}\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,B)$, we have $\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,E_{i})=0$ for all $i$. As $G$ weakly generates $\mathcal{D}^{\mathrm{cpt}}$, we have the result. ∎
In the proof of following 4.9, the homotopy colimit is used to approximate an object $X$ in a triangulated category with arbitrary direct sums by a set of objects in $\mathcal{D}^{\mathrm{cpt}}$ which is not necessarily finite.
Definition 4.7.
Let $\mathcal{D}$ be a triangulated category and $(X_{n},j_{n}:X_{n-1}\to X_{n})_{n\in\mathbb{Z}_{\geq 0}}$ be a system of $\mathcal{D}$. A homotopy colimit of the system $(X_{n})$ is defined as the objects fitting into the following triangle:
$\bigoplus_{n}X_{n}\stackrel{{\scriptstyle 1-j}}{{\longrightarrow}}\bigoplus_{n}X_{n}\to X\stackrel{{\scriptstyle[1]}}{{\to}}\Big{(}\bigoplus_{n}X_{n}\Big{)}[1],$
where
$1-j\coloneqq\begin{pmatrix}1&0&0&\cdots&&\\ -j_{1}&1&0&&&\\ 0&-j_{2}&1&&&\\ &\vdots&&\ddots&&\end{pmatrix}.$
Write this homotopy colimit $X$ by $\operatorname{hocolim}X_{n}$.
Lemma 4.8 ([Sta18, Tag 09SN]).
Let $\mathcal{D}$ be a triangulated category with direct sums and $\{E_{i}\}_{i\in I}\subset\operatorname{ob}(\mathcal{D}^{\mathrm{cpt}})$ objects such that generates $\mathcal{D}$. Then, for any $K\in\mathcal{D}$ there exists a sequence
$X_{1}\stackrel{{\scriptstyle j_{1}}}{{\to}}X_{2}\stackrel{{\scriptstyle j_{2}}}{{\to}}X_{3}\stackrel{{\scriptstyle j_{3}}}{{\to}}\cdots$
such that $X_{1}$ and each $\operatorname{Cone}(j_{i})$ are the direct sum of shifts of the $E_{i}$ and
$K=\operatorname{hocolim}X_{n}.$
Proposition 4.9.
Let $C$ be a curve and $G\in D^{b}(C)$. If $G$ weakly generates $D^{b}(C)$, $G$ classically generates $D^{b}(C)$.
Proof of 4.9.
Let $K\in D(\operatorname{\mathbf{Qcoh}}C)$ and assume that $\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(G,K)=0$. By the 4.8, there exists a sequence
$K_{1}\stackrel{{\scriptstyle j_{1}}}{{\to}}K_{2}\stackrel{{\scriptstyle j_{2}}}{{\to}}K_{3}\stackrel{{\scriptstyle j_{3}}}{{\to}}\cdots$
such that $K=\operatorname{hocolim}K_{n}$ and $\operatorname{Cone}(j_{i})=\bigoplus_{k}\mathcal{E}_{k}^{i}[-k]$,where $\mathcal{E}_{j}^{i}\in\operatorname{\mathbf{Coh}}(C)$. Here, for the convenience, write $\operatorname{Cone}(j_{0})=K_{1}$.
As any object in $D^{b}(C)$ splits, coherent sheaves $\mathcal{F}_{j}$ can rewrite $K_{i}$ as
$K_{i}\cong\bigoplus_{k}\mathcal{F}_{k}[-k].$
Thus, $K=\operatorname{Cone}(1-j)$ splits into the (infinite) direct sum of the shifts of the coherent sheaves. 4.6 claims that the $1-j$ is an isomorphism and therefore, $K\cong 0$. By 4.5, $G$ classically generates $D^{b}(C)$. ∎
Theorem 4.10.
Let $\Phi_{|\mathscr{O}_{C}(1)|}:C\hookrightarrow\mathbb{P}^{N}$ be a smooth projective curve with an embedding. Then, an Ulrich object exists on $(C,\mathscr{O}_{C}(1))$.
Proof of 4.10.
Note that $\mathscr{O}(1)\in D^{b}(C)$ cannot classically generate the whole derived cateogy $D^{b}(C)$. Indeed, as $\operatorname{rk}K_{0}(\langle\mathscr{O}(1)\rangle)=1$ by the stability of $\mathscr{O}(1)$, comparing the rank of the K-group of $\langle\mathscr{O}(1)\rangle$ and $D^{b}(C)$, we have
$1\displaystyle 2=\operatorname{rk}K_{0}(D^{b}(C))>\operatorname{rk}K_{0}(\langle\mathscr{O}_{C}(1)\rangle)=$ 1 .
Thus, $\mathscr{O}(1)$ cannot classically generate the whole derived category. Finally, 4.9 shows the claim. ∎
@@ 5.1. Preliminaries on Bridgeland Stability Conditions
Let $X$ be a surface, $D,H$ be an $\mathbb{R}$-divisors such that $H$ is ample. Define an abelian category by
@@
Let $X$ be a smooth projective surface over $\mathbb{C}$. Then, the divisorial stability condition $\sigma_{D,H}$ is geometric such that the phase of the skyscraper is 11 11. Conversely, if $\sigma\in\operatorname{Stab}(X)$ satisfies that $k(x)$ is $\sigma$-stable of phase 11 11 for any $x\in X$, then there exists a pair $(D,H)\in N^{1}(X)_{\mathbb{R}}\times\operatorname{Amp}(X)_{\mathbb{R}}$ such that $\sigma=\sigma_{D,H}$.
@@ 5.2.1. The Projective Plane
by 2.2. By the definition of $\mathcal{A}_{s,t}$, only one $n_{i}\neq 0$ if $E[k]\in\mathcal{A}_{s,t}$. ∎

### v2

@@ Abstract.
This paper gives a characterization of the Ulrich object similar to the one in [ES03]. As an application, we have provided a new approach to the Eisenbud-Schreyer question by using the notions of the generator of the derived category. We have also given an example of Ulrich objects that are not sheaves by the Yoneda extension.
@@
Derived category, Ulrich bundles, Algebraic curves
@@ 0. Introduction
Let $X$ be a smooth projective variety over $k=\overline{k}$ and $D^{b}(X)$ be the bounded derived category of coherent sheaves on $X$. A derived category, whose objects are bounded complexes of coherent sheaves, is a triangulated category constructed from an abelian category. The investigation of $D^{b}(X)$ as an invariant of a variety began with the work of Bondal and Orlov [BO01], where they showed that if $X$ has (anti-)ample canonical line bundle, then the equivalence $D^{b}(X)\cong D^{b}(Y)$ implies $X\cong Y$.
@@
Meanwhile, recent research has shown that the derived categories are also valuable to investigate the properties of sheaves. For instance, derived categorical methods are employed in the works of Ohno [Ohn20, Ohn22] and Fukuoka, Hara, and Ishikawa [FHI22, FHI23] to classify certain classes of special bundles.
@@ 0.3. Application to the Eisenbud-Schreyer’s Question
We present a new general proof of the existence of Ulrich bundles on an elliptic curve with an arbitrary embedding, based on the framework of the derived category. By B, known results, and a few discussions deduce the following corollary.
Theorem C (see 4.6).
For any elliptic curve with an embedding $(E,\mathscr{O}_{E}(1))$, there exists an Ulrich object.
Note that the existence of the Ulrich line bundle on a curve has already been established in [ES09], see also [Cos17] for a survey.
@@
The existence problem has been widely studied, and there are too many references to list them out here, but we mention some of them:
@@
The following Section 3 discusses the example of Ulrich objects. In particular, we introduce a special type of Ulrich object, called a Yoneda-type Ulrich object. The application of Ulrich object to A is given in Section 4. Especially, the proof of C is in Section 4.2.
@@
The conventions and notations used in this paper are listed below:
@@
  • all functors are derived unless otherwise stated. However, $\mathbf{L}$ and $\mathbf{R}$ are occasionally used to emphasize that the functor is actually derived.
@@ Acknowledgement.
The author is grateful to his advisor, Professor Yasunari Nagai, for his continuous encouragement and fruitful discussions. He would also like to thank Professor Yuki Mizuno for useful discussions and Professors Wahei Hara and Dmitri Pirozhkov for pointing out the mistake in Proposition 4.9 in the previous version of this paper. The author was supported by JST SPRING, Grant Number JPMJSP2128.
@@ 1.1. Preliminaries on Ulrich Bundles
$0\to\mathscr{O}_{\mathbb{P}^{n+1}}(-1)^{\oplus rd}\stackrel{{\scriptstyle M}}{{\longrightarrow}}\mathscr{O}_{\mathbb{P}^{n+1}}^{\oplus rd}\to\mathcal{E}\to 0.$
@@ 3.1. Yoneda Type Ulrich Objects
$0\to\mathcal{G}\to E_{k-1}\to\cdots\to E_{0}\to\mathcal{F}\to 0.$
@@
${0}$${\mathcal{G}}$${E_{k-1}}\displaystyle{\cdots}$${E_{1}}$${E_{0}}\displaystyle{\mathcal{F}}$${0}$${0}$${\mathcal{G}}$${G_{k-1}}\displaystyle{\cdots}$${G_{1}}$${G_{0}}\displaystyle{\mathcal{F}}$${0}$${0}$${\mathcal{G}}$${F_{k-1}}\displaystyle{\cdots}$${F_{1}}$${F_{0}}\displaystyle{\mathcal{F}}$${0.}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{\operatorname{id}}$
@@ 3.2. Explicit Examples
When the dimension of the variety is $1$, Ulrich objects are just a direct sum of Ulrich bundles from 2.2. Indeed, in this case, any object $E$ is formal, that is, $E$ can be written as
@@ 3.2.2. Ulrich Objects on Surfaces
Thus, giving an Ulrich object on a surface $S$ is equivalent to giving finitely many Ulrich sheaves $\mathcal{G}^{i}$ and their degree $2$ extensions.
Moreover, the Ulrichness of the object imposes a strong restriction on its Chern character. To see this, for example, let $X$ be a surface of Picard rank $1$. Let us write $\operatorname{Pic}(X)=\mathbb{Z}[H]$, $H^{2}=d$, and $K_{X}=i_{X}H$.
@@ 4.1. Generators of the Derived Category
$\displaystyle=\{E\in\mathcal{D}\mid\mathop{\mathbf{R}\mathrm{Hom}}\nolimits(X,E)=0,{}^{\forall}X\in\Omega\}.$
@@
4.2. Existence for Elliptic Curves
In this Section 4.2, we will show C.
Proposition 4.5.
Let $E$ be an elliptic curve and $G\in D^{b}(E)$. If $G$ weakly generates $D^{b}(E)$, $G$ classically generates $D^{b}(E)$.
Proof of 4.5.
We may assume $G$ is the direct sum of semistable sheaves $G_{i}$ $(1\leq i\leq m)$ from Atiyah’s classification. If $G$ is torsion or semistable, it is not a generator. Thus, we may assume $m\geq 2$ and $\mu(G_{1})\neq\mu(G_{2})$. Then $G$ must be a classical generator (see the proof of [BFK12, Lemma 6.7.] for instance). ∎
Theorem 4.6.
Let $\Phi_{|\mathscr{O}_{E}(1)|}:E\hookrightarrow\mathbb{P}^{N}$ be an elliptic curve with an embedding. Then, an Ulrich object exists on $(E,\mathscr{O}_{E}(1))$.
@@ 4.2. Existence for Elliptic Curves
Note that $\mathscr{O}(1)\in D^{b}(E)$ cannot classically generate the whole derived cateogy $D^{b}(E)$. Indeed, as $\operatorname{rk}K_{0}(\langle\mathscr{O}(1)\rangle)=1$ by the stability of $\mathscr{O}(1)$, comparing the rank of the K-group of $\langle\mathscr{O}(1)\rangle$ and $D^{b}(E)$, we have
$2=\operatorname{rk}K_{0}(D^{b}(E))>\operatorname{rk}K_{0}(\langle\mathscr{O}_{E}(1)\rangle)=1.$
Thus, $\mathscr{O}(1)$ cannot classically generate the whole derived category. Finally, 4.5 shows the claim. ∎
@@ 5.1. Preliminaries on Bridgeland Stability Conditions
Let $X$ be a surface, $D,H$ be an $\ RB$-divisor such that $H$ is ample. Define an abelian category by
@@
Let $X$ be a smooth projective surface over $\mathbb{C}$. Then, the divisorial stability condition $\sigma_{D,H}$ is geometric such that the phase of the skyscraper is $1$. Conversely, if $\sigma\in\operatorname{Stab}(X)$ satisfies that $k(x)$ is $\sigma$-stable of phase $1$ for any $x\in X$, then there exists a pair $(D,H)\in N^{1}(X)_{\mathbb{R}}\times\operatorname{Amp}(X)_{\mathbb{R}}$ such that $\sigma=\sigma_{D,H}$.
@@ 5.2.1. The Projective Plane
by $2.$ By the definition of $\mathcal{A}_{s,t}$, only one $n_{i}\neq 0$ if $E[k]\in\mathcal{A}_{s,t}$. ∎
@@ 5.2.3. Bridgeland Stability of Ulrich Objects on Surfaces
Lemma 5.9.
With the above settings,
$\displaystyle Z_{s,t}(E)$ $\displaystyle=\bigg{\{}-r\chi(\mathscr{O}_{X})+\frac{rd}{4}(i_{X}^{2}+3i_{X}+4)+\frac{rd}{2}\left(s^{2}-t^{2}+(i_{X}+3)s\right)\bigg{\}}$
$\displaystyle\quad+\bigg{\{}\frac{rd}{2}(2s+i_{X}+3)t\bigg{\}}\sqrt{-1}.$ (5.1)
Proof of 5.9.
We have the calculation using 3.6. ∎